OpenStudy (anonymous):

ric has a gift card to his favorite restaurant. Every Tuesday he has the same meal at the restaurant. What is the equation that represents Eric’s gift card balance? (y – 100) = (x – ) Eric’s Gift Card Weeks 2 4 6 8 Balance $125 $100 $75 $50

Directrix (directrix):

I used the points (6,75) and (8,50) to get the slope of this linear relationship. (50 - 75) /( 8 - 6) = -25/2 = - 12.5 From the given partial equation: (y – 100) = (x – ) I read off the chart that x = 4 is paired with 100. (y – 100) = (-25/2) * (x – 4) is the equation.----> @lllllllllllllllll You may prefer to call it (y – 100) = (-12.5) * (x – 4)